OpenStudy (anonymous):

Evaluate x/y for x=3/10 and y=4/5 a.7/50 b.7/15 c.3/8 d.12/50

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (anonymous):

Step 1: Formulate: \[\frac{ a }{ b }\] Step 2: Substitute \[\frac{ \frac{ 3 }{ 10 } }{ \frac{ 4 }{ 5 } }\] Step 3: Solve Use the outer and inner technique. Outer: 3 x 5 = 15 Inner: 10 x 4 = 40 So we have, 15/40 Step 4: Simplify 15/40 = 3/8 Answer: C. \[\frac{ 3 }{ 8 }\]

OpenStudy (anonymous):

Don't forget to click the "best response" if I helped. Thanks!

OpenStudy (anonymous):

Thank you again!

OpenStudy (anonymous):

You're welcome!